OpenStudy (anonymous):

solve m=2x+3y for y?

OpenStudy (anonymous):

y=m-2x/3

OpenStudy (anonymous):

Alright, I'll need your input to solve this, italian. You ready for this?

OpenStudy (anonymous):

y=(m-2x)/3

OpenStudy (anonymous):

Just giving them the answer won't help them learn lol You've gotta let them work through the answer. :P

OpenStudy (anonymous):

lol uber. sorry. shall delete my answer.

OpenStudy (anonymous):

Just giving them the answer won't help them learn lol You've gotta let them work through the answer. :P absolutely agree with this.

OpenStudy (anonymous):

Italian? Still here? We need to work through this if you wanna learn the steps. :P

OpenStudy (anonymous):

yeah im here

OpenStudy (anonymous):

Awesome! So which letter do we want by itself?

OpenStudy (anonymous):

y

OpenStudy (anonymous):

perfect! m=2x+3y

OpenStudy (anonymous):

First thing we want is for y to be positive. Is it?

OpenStudy (anonymous):

awesome! So this means we won't be subtracting y and that we'll be leaving it on the right side of the equals sign so that it stays positive. Now, we want y by itself. What is currently bothering it?

OpenStudy (anonymous):

Here's the equation again: m=2x+3y

OpenStudy (anonymous):

Yup! Now, here's the thing: for now we'll make believe that the 3 and y are buddies and don't want to separate.

OpenStudy (anonymous):

However... someone else is being really annoying. Who is that?

OpenStudy (anonymous):

Indeed. How do we make the 2x go away?

OpenStudy (anonymous):

Well, if we do that, we'd have to divide *everything* by 2x. That'd give us: m/2x = 3y/2x

OpenStudy (anonymous):

do we really want this?

OpenStudy (anonymous):

Note that 2y just got even closer to that 2x that was annoying him. :(

OpenStudy (anonymous):

no should you subtract?

OpenStudy (anonymous):

Perfect! So how will our equation look?

OpenStudy (anonymous):

m-2x=3y?

OpenStudy (anonymous):

Perfect!

OpenStudy (anonymous):

Now y can finally be by himself after one more step.

OpenStudy (anonymous):

then you divide both sides by 3 right?

OpenStudy (anonymous):

Heck ya! :D

OpenStudy (anonymous):

And what's our final answer? This step can be tricky.

OpenStudy (anonymous):

haha thanks a lot

OpenStudy (anonymous):

No problemo. Go ahead and add the finishing steps just to make sure we've got this. ^_^

OpenStudy (anonymous):

very close. Remember you have to divide everything by 3.

OpenStudy (anonymous):

There was one variable that you forgot to divide by 3.

OpenStudy (anonymous):

perfect!

OpenStudy (anonymous):

Alternatively, if you felt like it, m/3 - 2x/3 = y can also work :D

OpenStudy (anonymous):

Feel free to post another problem like this as a question to be on the safe side! More practice = more 100s! :P

OpenStudy (anonymous):

Of course. Feel free to make up a new separate question. :D

OpenStudy (anonymous):

You mean a different letter? It's already solved for t :P

OpenStudy (anonymous):

Ah ok.

OpenStudy (anonymous):

Well, let's start by trying to make r be less cramped. It's dangerous to work with things inside of parenthesis, so we should try to keep that for later.

OpenStudy (anonymous):

So... what CAN we do for now?

OpenStudy (anonymous):

Here's a hint with something similar: 5 = (3+7)/2

OpenStudy (anonymous):

Well, here's the thing: The example I showed is using totally different numbers, but is in the same general 'form' as your question, right?

OpenStudy (anonymous):

numbers in parentheses divided by a single number = another number, yes?

OpenStudy (anonymous):

In my example, the 7 is the equivalent to the P that we want to separate in your example. Does that part make sense?

OpenStudy (anonymous):

If it doesn't make sense, do let me know, and I'll explain it.

OpenStudy (anonymous):

here is the equation a different way \[t=p/6+3r/6\]

OpenStudy (anonymous):

To be honest, the original form of (p+3r) is to your advantage. However, it can also work here if you know the trick.

OpenStudy (anonymous):

Okie dokie.

OpenStudy (anonymous):

So going back to my 'similar example', we want to separate the 7, right? 5 = (3+7)/2 What can we do to my example to make the right side less messy?

OpenStudy (anonymous):

multiply by 2?

OpenStudy (anonymous):

perfect :D

OpenStudy (anonymous):

and what does that give us so far in my example?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

10=(3+7)

OpenStudy (anonymous):

Yup! Now that (3+7) is by itself, what can we get rid of?

OpenStudy (anonymous):

get rid of the ()

OpenStudy (anonymous):

Yup! and then how do you get the 7 by itself when you have 10 = 3+7?

OpenStudy (anonymous):

subtract the 3 from both sides

OpenStudy (anonymous):

Awesome. I lost your original problem. lol Could you repost it?

OpenStudy (anonymous):

The one with the P?

OpenStudy (anonymous):

t=(p+3r)/6 for r

OpenStudy (anonymous):

awesome. So what's the easiest way to make it look less ugly?

OpenStudy (anonymous):

multiply both sides by 6

OpenStudy (anonymous):

Indeed. What do we get?

OpenStudy (anonymous):

6t=p+3r

OpenStudy (anonymous):

yup! Now who is annoying our buddy r? (I think we were solving for r, right?)

OpenStudy (anonymous):

Hint: always remove items that can be moved easily to the other side first.

OpenStudy (anonymous):

yes and the friend p

OpenStudy (anonymous):

Yup. Make that p buzz off. We hate that guy.

OpenStudy (anonymous):

6t-p=3r? then divide both sides by 3 to get the answer?

OpenStudy (anonymous):

Yessssshhhhhh

OpenStudy (anonymous):

BUT, here's a tip: always check your answers.

OpenStudy (anonymous):

t=(p+3r)/6 for r

OpenStudy (anonymous):

let's plug in numbers to make this work. Let's make the parentheses equal 12 so that t can equal 2. That way we'll have 12/6 =2

OpenStudy (anonymous):

what we can do is... make r equal 2 and p equal 6. Make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

awesome. Could you go ahead and test it out for me, then? (Your answer, that is)

OpenStudy (anonymous):

here's what you came up with: 6t-p=3r turns into r = 2t - p/3

OpenStudy (anonymous):

so is 2 = 2(2)-6/3 ? Or 2 = 4-2?

OpenStudy (anonymous):

All is well? ^_^